OpenStudy (anonymous):

\[f(x)=\log_{(x-3)} (x-2)-\log_{(x-3)}(x^2+3x-10) \] What values of x will produce a defined function? (Domain)

jhonyy9 (jhonyy9):

do you know where is undefined the logarithm ?

jhonyy9 (jhonyy9):

and you need to know the property of log. like log(a) x - log(a) y = log(a) x/y hope these will help you sure

jhonyy9 (jhonyy9):

and the next step you can rewriting factorizing x^2 +3x -10 = (x -2)(x+5)

OpenStudy (anonymous):

I know that in logs for a function to be defined you must have: \[\log_{a} (b)\] a>0, a=/=1 b>0

OpenStudy (jmark):

f(x)=log_(x-3)(x-2)-log_(x-3)(x^2+3x-10) f(x)=[log(x-2)-log(x^2+3x-10)]/log(x-3) f(x)=[log(x-2)-log((x-2)(x+5))]/log(x-3) f(x)=-log(x+5)/log(x-3) the domain {x belongs to R:3<x<4 or x>4}

OpenStudy (anonymous):